OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -2) and (9, -2)? 3 0 -3 undefined

OpenStudy (anonymous):

I just gave you the slope formula, try and use it.

OpenStudy (anonymous):

\[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (anonymous):

-2-(-2)/9-3 0/3 0

OpenStudy (dan815):

yep, because there is no change in y, it must be a horizontal line, so slope 0

OpenStudy (dan815):



OpenStudy (anonymous):

can you help with this one too The slope of the graph of a direct variation function is 4. What is the equation of the function? y =1/4 x y = 4x + 1 y = -4x y = 4x

OpenStudy (dan815):

slope = change in y/change in x

OpenStudy (tkhunny):

No. -2-(-2)/9-3 = -2 + 2/9 - 3 = -16/9 - 3 = -43/9 Now, [-2-(-2)]/[9-3] = 0/6 = 0 The extra parentheses are NOT optional.

OpenStudy (dan815):

y=mx+b <--- equation of a line, m is the slope and b is the yintercept

OpenStudy (dan815):

direct variation is given by only y=mx <---- m is the slope what do you think m should be?